# Top-of-atmosphere radiation fields over the last millennium reconstructed from proxies

### v1

@@ Abstract
Earth’s energy imbalance at the top of the atmosphere is a key climate system metric, but its variability is poorly constrained by the short observational record and large uncertainty in coupled climate models. While existing ocean heat content reconstructions offer a longer perspective, they cannot separate the contributions of shortwave and longwave radiation, obscuring the underlying processes. We extend the energy budget record by reconstructing the top-of-atmosphere radiation and related surface variables over the last millennium (850–2000 CE). Our method combines proxy data and model dynamics using seasonal, gridded data assimilation, exploiting the covariance of radiation with proxies sensitive to surface temperatures. The method validates well in a pseudoproxy experiment and against instrumental observations. We find that a last-millennium cooling trend coincides with heat loss that gradually slowed, although there are intermittent multidecadal periods of energy gain. The cooling trend is associated with a southwestward shift of Indo–Pacific convection and growth of sea ice, with seasonal sea ice trends following orbital-scale changes in polar insolation. We also find that the upper-ocean heat content following large volcanic eruptions does not begin to recover until 5–10 years later, suggesting the initiation of the Little Ice Age by decadally-paced eruptions in the early 1100s and late 1200s. Indeed, the latter period marks the decade of largest energy loss over the last millennium. Our reconstruction reveals that the energy imbalance for all 20-yr periods after 1980 is unprecedented in the pre-industrial period.
@@ 1 Introduction
Earth’s energy imbalance (EEI) at the top of the atmosphere (TOA) is a fundamental climate system metric, governing the total energy inventory and constraining global temperatures, the hydrological cycle, sea levels, and ice cover (von Schuckmann et al., 2016). Much of the EEI is forced, over the last millennium mostly by volcanic eruptions, solar irradiance, and changes in Earth’s orbit, and more recently by anthropogenic greenhouse gases and aerosols. In addition, there is internal, unforced variability in the energy budget at all timescales. Seasonal variability is primarily due to extratropical storms and the Madden–Julian oscillation, while the El Niño–Southern Oscillation (ENSO) explains most interannual variability (Trenberth et al., 2015). However, decadal to centennial variability in the energy budget remains poorly understood (Wills et al., 2021; Trenberth et al., 2014). This complicates, for example, the interpretation of recent albedo trends (Goessling et al., 2025; Mauritsen et al., 2025; Hodnebrog et al., 2024; Raghuraman et al., 2021) and the warming hiatus (Medhaug et al., 2017; Xie et al., 2016; Trenberth and Fasullo, 2013), both of which are not reproduced well by coupled climate models. A better understanding of Earth’s energy budget would also help to attribute historical climate change (Lean, 2018) and constrain future warming (Sherwood et al., 2020).
@@
Our understanding of EEI variability is limited by the short satellite observational record of around 25 years (Loeb et al., 2024), which prohibits the investigation of low-frequency variability and the response to rare, episodic forcings. Additionally, it is difficult to disentangle forced and internal variability of Earth’s energy budget due to the strong anthropogenic greenhouse gas forcing and the uncertain aerosol forcing during the observed period. Therefore, coupled climate models have been used to investigate aspects like radiative feedbacks (Sherwood et al., 2020) and decadal variability (Palmer and McNeall, 2014; Brown et al., 2014; Zhou et al., 2016a), more recently through coordinated experiments (Schmidt et al., 2023; Webb et al., 2017). However, even models of the latest generation produce a large spread in TOA radiation, on the order of 5–20 W m-2 for most components (Wild, 2020), and systematically underestimate the radiative response (Olonscheck and Rugenstein, 2024). In addition, simulations of the last millennium disagree with proxy evidence on global temperature trends, their realization of internal variability often differs from the true climate history, and specified forcings may not be accurate (Lean, 2018; Lücke et al., 2023; Fyfe et al., 2021). Other studies have extended the energy budget record by reconstructing the ocean heat content (OHC). For example, Zanna et al. (2019) and Wu et al. (2025) derive the global-mean OHC over the historical period, while Gebbie and Huybers (2019) reconstruct the gridded OHC over the Common Era. Further OHC reconstructions from the Last Glacial Maximum to present are reviewed in Gebbie (2021). On interannual to decadal timescales, the OHC is a more reliable indicator of global change than surface temperatures (Palmer and McNeall, 2014; Allison et al., 2020). However, the ocean perspective, particularly in the global mean, obscures the atmospheric processes that cause heat content changes, and the role of clouds and sea ice in mediating them. Rather, a top-of-atmosphere perspective that separates shortwave (SW) and longwave (LW) components is needed to improve the process understanding of energy budget variability.
@@
Here, we reconstruct Earth’s energy budget over the last millennium (850–2000 CE) from proxies. The reconstruction is seasonally and spatially resolved, and partitions the TOA radiation into SW and LW components. We use data assimilation, which combines proxy observations with model dynamics, and has been used successfully to reconstruct past climates (e.g., Goosse et al., 2010; Perkins and Hakim, 2021; Meng et al., 2025; Hakim et al., 2016; Valler et al., 2024; Judd et al., 2024; Cooper et al., 2025). Our method exploits the covariance of TOA radiation and OHC with surface temperatures, to which the proxies are sensitive. For example, if a proxy in the Pacific Warm Pool indicates warmer sea surface temperatures (SSTs), the atmospheric response there is to enhance deep convection, which mediates an increase in reflected SW and a decrease in outgoing LW radiation (Fig. 1; e.g., Dong et al., 2019). Similarly, changes in temperature gradients can affect the overturning atmospheric circulation with implications for subtropical low clouds, which have a strong SW effect. Besides clouds, snow/ice cover and solar surface heating play a role, and establish a temperature–radiation relationship over much of the globe (Trenberth et al., 2015). In addition to this physical underpinning, Loeb et al. (2020) demonstrates that the current generation of atmosphere-only climate models can skillfully simulate TOA radiation when SST and sea ice are prescribed; both of these variables can be reconstructed from proxies.
@@
Figure 1: Annual-mean correlation between the SST at a point in the Niño 3.4 region (yellow pentagon; e.g., location of a coral proxy) and the TOA (a) reflected SW and (b) outgoing LW radiation globally. The correlation is from last-millennium simulations over 850–1850, averaged over four models.
@@
We take two perspectives on the TOA energy balance. The first perspective involves a partition of the EEI into insolation $S$, reflected SW radiation (RSR), and outgoing LW radiation (OLR):
@@
By our chosen sign convention, EEI is positive downwelling (energy gain), while RSR and OLR are positive upwelling (energy loss). Separating the SW and LW radiation fields allows us to infer the physical processes governing the EEI, particularly clouds. The reconstructed partitioning is derived from emulators of last-millennium climate simulations and is therefore not aimed at long-term trends during periods of strong anthropogenic forcing from greenhouse gases and aerosols (i.e., the 1900s).
The second perspective on the TOA energy balance emphasizes the relation to the global-mean surface temperature $T$ in a one-layer energy balance model (Geoffroy et al., 2013):
@@
where EEI is the global-mean TOA energy imbalance, $C$ is the heat capacity of the upper ocean, and $\gamma$ is the heat uptake coefficient of the deep ocean. On short timescales the deep-ocean heat uptake is relatively small, so we can approximately compare OHC and EEI by differentiation or integration. A third perspective on EEI involves the forcing–feedback framework (e.g., Sherwood et al., 2015). Both forcings and feedbacks correlate with temperature-sensitive proxies, forcings through their cumulative effect on heat content, feedbacks through a temperature-mediated response. However, we can only reconstruct their sum using our method. Separating forcings and feedbacks requires climate model simulations and is beyond the scope of this study.
The remainder of the paper is organized as followed. We first describe the data assimilation method (Section 2). After validating this method in a pseudoproxy experiment and the real-proxy reconstruction against instrumental datasets (Section 3), we present our reconstructed energy budget over the last millennium (Section 4), then illustrate applications to the pre-industrial energy budget variability and the response to volcanic eruptions (Section 5).
@@ 2 Methods
We use online paleoclimate data assimilation (DA) to reconstruct gridded climate fields at seasonal resolution. DA combines proxy information with forecasts from a model in a statistically optimal way (e.g., Kalnay et al., 2024). Our ensemble DA system comprises linear inverse models (LIMs), an ensemble Kalman filter (EnKF), and proxy system models (PSMs). The methodology for our seasonal-resolution DA is similar to Meng et al. (2025) and based on the prior work from Perkins and Hakim (2021) and Hakim et al. (2016).
Ensemble DA represents the distribution of the state in each season using samples, each of which is equally likely. The procedure starts with a climatological prior. During the update step, the EnKF updates the prior ensemble with available proxy data to obtain the posterior ensemble. During the forecast step, the LIM propagates all posterior ensemble members to the next season, where the forecast serves as prior for another update with available proxies. The posteriors from this repeated forecast–update cycle comprise the reconstruction.
@@ 2.1 Forecasts using linear inverse models
We use LIMs as efficient emulators of the dynamics and statistics of coupled climate models. Using such emulators allows us to run large ensembles and incorporate the dynamics of different climate models. The LIM dynamics have the form
@@
where $\mathbf{x}$ is the state vector, $\bm{\mathsf{L}}$ is a constant matrix representing the linear dynamics, $\bm{\mathsf{S}}$ is the noise amplitude matrix, and $\bm{\eta}$ is a vector of Gaussian white noise with unit variance. Together, $\bm{\mathsf{L}}$ and $\bm{\mathsf{S}}$ encode, for example, how surface temperatures and TOA radiation are related. The full equations for determining the matrices and for ensemble forecasts are presented in the Supplemental Information.
@@
The state vector $\mathbf{x}$ represents seasonal anomalies of atmospheric and oceanic climate variables. We include 2-m surface air temperature (SAT), sea surface temperature (SST), TOA energy imbalance (EEI), TOA reflected SW radiation (RSR), TOA outgoing LW radiation (OLR), ocean heat content of the upper 300 m (OHC300), Arctic sea ice concentration (SIC), and Antarctic SIC. We integrate the OHC over 300 m, which is the depth required to close the seasonal energy budget (Johnson et al., 2023) and is sufficiently deep to remove surface noise (Allison et al., 2020). The OHC also imparts memory to the LIM forecasts since anomalies persist longer than in atmospheric fields.
@@
The training data are regridded to the same $\times 2$° grid, converted to anomalies relative to their own 1961–1990 climatology, then linearly detrended by season to remove the orbital precession signal. Any trend in the reconstruction is therefore due to proxies. The full gridded state has significant spatial correlations at seasonal resolution, motivating a large dimensional reduction for the LIM. We accomplish this reduction by using empirical orthogonal functions (EOF) on the area-weighted gridded data, and constructing the state vector from the truncated principal components (PCs):
$\mathbf{x}=[\tilde{\mathbf{x}}_{20}^{\text{SAT}};\tilde{\mathbf{x}}_{20}^{\text{SST}};\tilde{\mathbf{x}}_{30}^{\text{EEI}};\tilde{\mathbf{x}}_{15}^{\text{RSR}};\tilde{\mathbf{x}}_{10}^{\text{OLR}};\tilde{\mathbf{x}}_{15}^{\text{OHC300}};\tilde{\mathbf{x}}_{10}^{\text{SICn}};\tilde{\mathbf{x}}_{10}^{\text{SICs}}],$
@@
where semicolons denote vertical stacking, and $\tilde{\mathbf{x}}_{n}^{\text{var}}$ represents the leading $n$ PCs for variable var, standardized by the square root of the retained variance after truncation. SICn/SICs refer to Arctic/Antarctic SIC. The truncation ranks $n$ were chosen subjectively based on the cumulative variance explained, trading off reconstruction fidelity and degrees of freedom required for LIM training. We use the same number of PCs for each model prior, despite differences in the explained variance (Fig. S1). This ensures a fair comparison since the number of degrees of freedom is comparable.
@@
The dynamical and stochastic components of the LIM are determined from climate model simulations. We train LIMs on the CMIP6 last-millennium simulations (past1000 and past2k; Jungclaus et al., 2017) from the following models: MPI-ESM1-2-LR, CESM2-WACCM-FV2, and MRI-ESM2-0. These are the “model priors”, hereafter referred to as MPI, CESM, and MRI (Mauritsen et al., 2019; Danabasoglu et al., 2020; Yukimoto et al., 2019). We exclude MIROC-ES2L (MIROC) from LIM training due to its Antarctic sea ice low bias (Hajima et al., 2020). Last-millennium simulations from other models exist but either lack needed variables, like SIC, have issues with the implementation of forcings, or have not been published. Previous reconstructions like Perkins and Hakim (2021) and Meng et al. (2025) train LIMs on CMIP5 models. However, only since the CMIP6 generation are these models able to faithfully simulate TOA radiation, mainly due to better representations of the low-cloud SW effect (Loeb et al., 2020).
Each LIM is used separately in the DA algorithm to assimilate proxies, yielding three reconstructions, each containing 400 members (details provided below). Multi-model ensembles have proven effective in reducing reconstruction error, particularly in regions distant from proxies (Parsons et al., 2021).

### v2

# Top-of-atmosphere radiation over the last millennium reconstructed from proxies

@@ Abstract
Earth’s energy imbalance at the top of the atmosphere is a key climate system metric, but its natural variability is poorly constrained by the short observational record and large uncertainty in coupled climate models. While existing ocean heat content reconstructions offer a longer perspective, they cannot separate the contributions of shortwave and longwave radiation, obscuring the underlying processes. We extend the energy budget record into the pre-industrial period by reconstructing the top-of-atmosphere radiation and related surface variables over the last millennium (850–2000 CE) by using data assimilation to combine proxy data and dynamics from a coupled climate emulator. Validation reveals skill in the reconstructed radiation fields, especially in the tropics. Results show a familiar last-millennium cooling trend, which coincides with persistent heat loss and a reduction in upper-ocean heat content. The cooling trend differs by season and latitude, and is associated with radiative anomalies suggestive of an eastward shift in Indo–Pacific convection. Following large volcanic eruptions, ocean heat content anomalies persist for 10–20 years on average, supporting previous evidence that the cooling trend was forced by decadally-paced eruptions. The reconstruction also reveals that the current rate of energy gain is unprecedented relative to the period before 1850.
@@ 1 Introduction
Earth’s energy imbalance (EEI) at the top of the atmosphere (TOA) is a fundamental climate system metric, governing Earth’s total heat content and constraining global temperatures, the hydrological cycle, sea levels, and ice cover (von Schuckmann et al., 2016). Much of the EEI is forced, over the last millennium mostly by volcanic eruptions, solar irradiance, and changes in Earth’s orbit, and more recently by anthropogenic greenhouse gases and aerosols. In addition, there is internal, unforced variability in the energy budget at all timescales. Seasonal variability is primarily due to extratropical storms and the Madden–Julian oscillation, while the El Niño–Southern Oscillation (ENSO) explains most interannual variability (Trenberth et al., 2015). However, decadal to centennial variability in the energy budget remains poorly understood (Wills et al., 2021; Trenberth et al., 2014). This complicates, for example, the interpretation of recent albedo trends (Goessling et al., 2025; Mauritsen et al., 2025; Hodnebrog et al., 2024) and the warming hiatus (Trenberth and Fasullo, 2013; Xie et al., 2016), both of which are not reproduced well by coupled climate models (Olonscheck and Rugenstein, 2024; Raghuraman et al., 2021; Medhaug et al., 2017; Fyfe et al., 2013). A better understanding of natural variability in Earth’s energy budget is also essential to isolating the contribution of anthropogenic forcing in observations, which informs the attribution of historical climate change (Lean, 2018) and helps narrow uncertainty in climate sensitivity (Sherwood et al., 2020).
@@
Our understanding of EEI variability is limited by the short satellite observational record of around 25 years (Loeb et al., 2024), which prohibits the investigation of low-frequency variability and the response to rare, episodic forcings. Even during this well-observed satellite period, discrepancies and uncertainties between EEI estimates are large (Hakuba et al., 2024). Additionally, it is difficult to disentangle forced and internal variability of Earth’s energy budget due to the strong anthropogenic greenhouse gas forcing and the uncertain aerosol forcing during the observed period. Therefore, climate models have been used to investigate aspects of the energy budget like radiative feedbacks (Sherwood et al., 2020) and decadal variability (Palmer and McNeall, 2014; Brown et al., 2014; Zhou et al., 2016), more recently through coordinated experiments like CFMIP and CERESMIP (Webb et al., 2017; Schmidt et al., 2023). However, even coupled models of the CMIP6 generation (Eyring et al., 2016) produce a large spread in TOA radiation, on the order of $5$ to $20\,\mathrm{W\,m^{-2}}$ for most components (Wild, 2020), and systematically underestimate the radiative response to surface warming (Olonscheck and Rugenstein, 2024). In addition, the specified forcings and simulated internal variability may not be accurate (Lean, 2018; Lücke et al., 2023; Fyfe et al., 2021). Other studies have extended the energy budget record by reconstructing the ocean heat content (OHC), which is storing much of the energy gained in recent decades (von Schuckmann et al., 2023). For example, Zanna et al. (2019) and Wu et al. (2025) derive the global-mean OHC over the historical period, while Gebbie and Huybers (2019) reconstruct the gridded OHC over the Common Era. Further OHC reconstructions from the Last Glacial Maximum to present are reviewed in Gebbie (2021). The OHC provides a valuable perspective since it is a more reliable indicator of global change than surface temperatures on interannual to decadal timescales (Palmer and McNeall, 2014; Allison et al., 2020). However, the OHC perspective on the energy budget, particularly in the global mean, obscures the atmospheric processes that cause heat content changes, and the role of clouds and sea ice in mediating them. Rather, a top-of-atmosphere perspective that separates shortwave (SW) and longwave (LW) components is needed to improve the process understanding of energy budget variability.
@@
Here, we reconstruct Earth’s energy budget over the last millennium (850–2000 CE) from paleoclimate proxies. The reconstruction is seasonally and spatially resolved, and partitions the TOA radiation into SW and LW components. We use data assimilation, which combines proxy observations with model dynamics, and has been used successfully to reconstruct past climates (Goosse et al., 2010; Bhend et al., 2012; Steiger et al., 2014, 2018; Perkins and Hakim, 2021; Hakim et al., 2016; Franke et al., 2017; Osman et al., 2021; Erb et al., 2022; Valler et al., 2022, 2024; Judd et al., 2024; Meng et al., 2025; Cooper et al., 2025, e.g.,; see Smerdon et al., 2023 for an overview). Our method exploits the covariance of TOA radiation and OHC with surface temperatures, to which the proxies are sensitive. For example, if a proxy in the Pacific Warm Pool indicates warmer sea surface temperatures (SSTs), the atmospheric response there is to enhance deep convection, which mediates an increase in reflected SW and a decrease in outgoing LW radiation (Fig. 1; e.g., Dong et al., 2019). Similarly, changes in temperature gradients can affect the overturning atmospheric circulation and tropospheric stability, which affects SW radiation through changes in subtropical low clouds (e.g., Mackie et al., 2025; Ceppi and Gregory, 2017; Van Loon and Rugenstein, 2025). Besides clouds, snow/ice cover and solar surface heating play a role, and establish a temperature–radiation relationship over much of the globe (Trenberth et al., 2015). In addition to this physical underpinning, Loeb et al. (2020) demonstrate that the current generation of atmosphere-only climate models can skillfully simulate TOA radiation when SST and sea ice are prescribed; both of these variables can be reconstructed from proxies.
@@
Figure 1: Annual-mean correlation between the SST at a point in the Niño 3.4 region (yellow pentagon; e.g., location of a hypothetical coral proxy) and the TOA (a) reflected SW and (b) outgoing LW radiation globally. Correlations are from last-millennium simulations over 850–1850, averaged over five models using the Fisher z-transformation.
@@
We take three perspectives on the TOA energy balance. The first perspective involves a partition of the EEI into insolation $S$, reflected SW radiation (RSR), and outgoing LW radiation (OLR):
@@
By our chosen sign convention, EEI is positive downwelling (energy gain), while RSR and OLR are positive upwelling (energy loss). Separating the SW and LW radiation fields allows us to infer the physical processes governing the EEI, such as clouds.
The second perspective on the TOA energy balance emphasizes the relation to the global-mean surface temperature $T$ in a one-layer energy balance model (Geoffroy et al., 2013; Gregory et al., 2016):
@@
where EEI is the global-mean TOA energy imbalance, $C$ is the heat capacity of the upper ocean, and $\gamma$ is the heat uptake coefficient of the deep ocean. The product $CT$ is the upper-ocean heat content, which we also reconstruct. On seasonal timescales, such as for the initial response to volcanic cooling, before the upper ocean has adjusted significantly, the deep-ocean heat uptake $\gamma T$ is relatively small. We can then approximately compare OHC and EEI by differentiation or integration. On interannual and longer timescales, the assumption of $\gamma T\ll C(dT/dt)$ breaks down (e.g., Jeevanjee et al., 2025).
A third perspective on EEI involves the forcing–feedback framework (e.g., Sherwood et al., 2015):
$\text{EEI}=R(T)+F_{\text{nat}}+F_{\text{anthro}},$ (2)
where $R(T)$ is the radiative response (also called feedback), and the forcing is separated into natural ($F_{\text{nat}}$; i.e., volcanic and solar) and anthropogenic ($F_{\text{anthro}}$) components. Both forcing and response correlate with temperature-sensitive proxies. Forcing relates to near-surface temperature through its cumulative effect on ocean heat content, and the response through temperature-mediated processes such as Planck and cloud feedbacks (Forster et al., 2021). Thus, the energy imbalance both causes and results from temperature changes. Our method, targeting the pre-industrial last millennium, restricts the reconstruction to the sum of $R(T)+F_{\text{nat}}$, without anthropogenic forcing.
The remainder of the paper is organized as follows. We first describe the data assimilation method (Section 2). After validating this method in pseudoproxy experiments and against instrumental datasets (Section 3), we present our reconstructed energy budget over the last millennium (Section 4), then illustrate applications to the natural energy budget variability and the response to volcanic eruptions (Section 5). Section 6 presents the conclusions.
@@ 2 Methods
We use online paleoclimate data assimilation (DA) to reconstruct gridded climate fields at seasonal resolution. DA combines proxy information with model forecasts in a statistically optimal way (e.g., Kalnay et al., 2024). Our DA system comprises three interacting components: (1) linear inverse models (LIMs; Penland and Sardeshmukh, 1995), trained to emulate the last-millennium climate dynamics of CMIP6 models, are used for forecasting; (2) proxy system models (PSMs; Dee et al., 2016; Evans et al., 2013, e.g.,) map from the climate variables predicted by the model to proxy values (e.g., from local SST to coral $\delta^{18}$O); and (3) an ensemble Kalman filter (EnKF; Evensen, 1994) that blends proxy data with model forecasts weighted by their uncertainties, spreading information across space and climate variables. Together, these components provide an estimate of the evolving state of the climate system by continuously updating a model forecast, which maintains memory of the climate from past proxies, with proxy observations during the forecast period.
Specifically, the interaction of the three components proceeds through a forecast–update cycle. A LIM produces a forecast “prior” given an initial state. The first initial state is a random draw from the model climatology, whereas subsequent initial states come from the cycle’s update step described next. The PSMs map this prior state to the expected proxy values, and the EnKF compares these expected values against the actual proxy observations, updating the prior state to produce a posterior estimate. This completes the cycle, which continues in time by using the posterior state as the new initial state to forecast the next season. The methodology for our seasonal-resolution DA is similar to Meng et al. (2025) and based on the prior work by Perkins and Hakim (2021) and Hakim et al. (2016).
@@ 2.1 Forecasts using linear inverse models
We use LIMs as efficient emulators of the dynamics and statistics of the last-millennium climate as simulated in five CMIP6 models. Using such emulators allows us to run large ensembles and to incorporate the dynamics of multiple climate models. The LIM dynamics have the form (Penland and Sardeshmukh, 1995; Penland, 1996)
@@
where $\mathbf{x}$ is the state vector, $\bm{\mathsf{L}}$ is a matrix representing the linear dynamics, $\bm{\mathsf{S}}$ is the noise amplitude matrix, and $\bm{\eta}$ is a vector of independent Gaussian white noise with unit variance. Together, $\bm{\mathsf{L}}$ and $\bm{\mathsf{S}}$ encode, for example, how surface temperatures and TOA radiation are related. During LIM training, the deterministic dynamics $\bm{\mathsf{L}}$ are derived from a lag-time linear regression, and the noise amplitude $\bm{\mathsf{S}}$ follows from a fluctuation–dissipation relation such that the long-term covariance statistics of the training data are reproduced. During forecasting from one season to the next, each ensemble member starts from a different initial condition and uses a different realization of $\bm{\eta}$. The full equations for determining the LIM matrices and for ensemble forecasts are presented in the Supplemental Information Text S1.
@@
The state vector $\mathbf{x}$ represents seasonal anomalies of atmospheric and oceanic climate variables. During LIM training, the state vector is constructed from last-millennium simulations. During DA, the goal is to estimate the posterior $\mathbf{x}$ by blending the LIM forecast (i.e., the prior $\mathbf{x}$) and proxy information. We include 2-m surface air temperature (SAT), sea surface temperature (SST), TOA energy imbalance (EEI), TOA reflected SW radiation (RSR), TOA outgoing LW radiation (OLR), ocean heat content of the upper 300 m (OHC300), Arctic sea-ice concentration (SIC), and Antarctic SIC. We integrate the OHC over 300 m, which is the depth required to close the seasonal energy budget (Johnson et al., 2023) and is sufficiently deep to remove surface noise (Allison et al., 2020). The OHC also imparts memory to the LIM forecasts since anomalies persist longer than in atmospheric fields.
@@
The LIM training data, consisting of last-millennium simulations over 850–1850, are regridded to the same $\times 2$° grid, converted to anomalies relative to their own 850–1850 climatology, then linearly detrended by season to remove the orbital precession signal. Any low-frequency trend in the reconstruction is therefore driven by the assimilated proxies.
Because the physical gridded state exhibits significant spatial correlation at seasonal resolution, we perform a large dimensionality reduction prior to LIM training. We accomplish this by first subtracting the global means from the gridded state, then calculating the empirical orthogonal functions (EOF) of the area-weighted gridded residuals. Explicitly separating the global means preserves their variance, which is particularly beneficial for spatially noisy fields such as RSR. We then construct the state vector from the global means and truncated principal components (PCs):
$\mathbf{x}=[\tilde{\mathbf{x}}_{1+20}^{\text{SAT}};\tilde{\mathbf{x}}_{1+20}^{\text{SST}};\tilde{\mathbf{x}}_{1+15}^{\text{EEI}};\tilde{\mathbf{x}}_{1+15}^{\text{RSR}};\tilde{\mathbf{x}}_{1+10}^{\text{OLR}};\tilde{\mathbf{x}}_{1+15}^{\text{OHC300}};\tilde{\mathbf{x}}_{10}^{\text{SICn}};\tilde{\mathbf{x}}_{10}^{\text{SICs}}],$
@@
where semicolons denote vertical stacking, and $\tilde{\mathbf{x}}_{1+n}^{\text{var}}$ represents the global mean, followed by the leading $n$ PCs for variable var. For Arctic and Antarctic sea ice (SICn and SICs), we do not separate the global mean and use only the leading 10 PCs. The global mean is standardized by its temporal standard deviation, and the PCs are standardized by the square root of their retained variance after truncation. The truncation ranks $n$ were selected subjectively based on the cumulative variance explained, trading off reconstruction fidelity and degrees of freedom required for LIM training. To ensure a fair comparison across model priors, we use the identical number of PCs for each, despite variations in the actual variance explained (Fig. S1). Mapping this EOF-space state vector $\mathbf{x}$ back to the physical gridded state amounts to a simple linear transformation.
@@
We train LIMs on the CMIP6 last-millennium simulations (“past1000” and “past2k”; Jungclaus et al., 2017) from the following five models: MPI-ESM1-2-LR, CESM2-WACCM-FV2, MRI-ESM2-0, EC-Earth3-Veg-LR, and MIROC-ES2L (Mauritsen et al., 2019; Danabasoglu et al., 2020; Yukimoto et al., 2019; Döscher et al., 2022; Hajima et al., 2020). These are the “model priors,” hereafter referred to as MPI, CESM, MRI, EC-Earth, and MIROC. Last-millennium simulations from other models exist but either lack necessary variables, like SIC, or have not been published. Previous reconstructions like Perkins and Hakim (2021) and Meng et al. (2025) train LIMs on CMIP5 models. However, only since the CMIP6 generation are these models able to faithfully simulate TOA radiation, mainly due to better representations of the low-cloud SW effect (Loeb et al., 2020). We exclude Antarctic sea ice from the MIROC-based LIM during our analysis due to its Antarctic SIC low bias (Hajima et al., 2020).
LIM forecast skill from season to season is important to propagate proxy information through time and to provide accurate priors. Forecast tests show skillful forecast correlations for the global means ($r>0.8$ for all fields at one season lead time; Fig. S2). The LIM tends to be slightly underdispersive, with ensemble spread underestimating forecast error in the ensemble mean (not shown). Spatial skill varies from $r\approx 0.5$ for SAT and OHC300, to $r\approx 0.3$ for radiation fields, with locally better skill in the tropical Pacific. Spatial correlation is primarily lost due to EOF truncation rather than lack of LIM skill, as evident from the reduced correlations of the initial condition. Conversely, the correlations for the global mean are close to unity, highlighting the value of handling it separately in the state vector.
Training on last-millennium simulations means that our reconstruction is most representative for the 850–1850 period. During the historical period (1850 onwards), the energy imbalance is dominated by anthropogenic forcing, for which the relationship with surface temperatures is different than the one learned during the last millennium. Our reconstructed EEI therefore only includes the response and natural forcing, even during the historical period.